OpenStudy (anonymous):

A number is chosen at random from all two-digit numbers. Find the probability that the number is (a) less than 40; (b) a multiple of 5; (c) not a multiple of 5.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

how many two digit numbers are there ?

ganeshie8 (ganeshie8):

and, how many two digit numbers are less than 40 ?

ganeshie8 (ganeshie8):

two digit numbers :- 10, 11, 12.... 99 can u count them ha ? :)

OpenStudy (anonymous):

less than 40, does it means 40 also?

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

less than or equal to 40, means 40 also

ganeshie8 (ganeshie8):

less than 40 : 10, 11, .... 38, 39

OpenStudy (anonymous):

i forgot the formula of calculating the number of terms..

ganeshie8 (ganeshie8):

there is no formula, just count them :)

OpenStudy (anonymous):

i remember it has a formula... ok. 90 terms and 30 terms?

ganeshie8 (ganeshie8):

thats right ! for probability, take the ratio fo them

OpenStudy (anonymous):

ratio? 3:1 ?

ganeshie8 (ganeshie8):

read the question again you are randomly pickign 2 digit numbers, and want to find the probability for it to be less than 40

OpenStudy (anonymous):

1/3

ganeshie8 (ganeshie8):

number of ways it can be less than 40 = 30 total number of ways = 90 so probability = number of ways in favor / total number of ways = 30/90 = 1/3

ganeshie8 (ganeshie8):

yes :) try part b

OpenStudy (anonymous):

10, 15, 20, 25, 30, 35, 40, 45, 50, 55, 60, 65, 70, 75, 80, 85, 90, 95 18 terms

OpenStudy (anonymous):

18/90 = 1/5

ganeshie8 (ganeshie8):

correct !

OpenStudy (anonymous):

not a multiple of 5 = 1- 1/5 = 4/5 ?

ganeshie8 (ganeshie8):

thats right !!

OpenStudy (anonymous):

thank you so much!

ganeshie8 (ganeshie8):

np :)